### v1

# SynQuE: Estimating Synthetic Dataset Quality Without Annotations

Abstract
We introduce and formalize the Synthetic Dataset Quality Estimation (SynQuE) problem: ranking synthetic datasets by their expected real-world task performance using only limited unannotated real data. This addresses a critical and open challenge where data is scarce due to collection costs or privacy constraints. We establish the first comprehensive benchmarks for this problem by introducing and evaluating proxy metrics that choose synthetic data for training to maximize task performance on real data. We introduce the first proxy metrics for SynQuE by adapting distribution and diversity-based distance measures to our context via embedding models. To address the shortcomings of these metrics on complex planning tasks, we propose Lens, a novel proxy that leverages large language model reasoning. Our results show that SynQuE proxies correlate with real task performance across diverse tasks, including sentiment analysis, Text2SQL, web navigation, and image classification, with Lens consistently outperforming others on complex tasks by capturing nuanced characteristics. For instance, on text-to-SQL parsing, training on the top-3 synthetic datasets selected via SynQuE proxies can raise accuracy from 30.4% to 38.4 (+8.1)% on average compared to selecting data indiscriminately. This work establishes SynQuE as a practical framework for synthetic data selection under real-data scarcity and motivates future research on foundation model-based data characterization and fine-grained data selection.
1 Introduction
Data scarcity hinders effective machine learning, especially for tasks requiring specialized expertise like autonomous navigation or natural language interfaces, where data collection is costly and slow (xie_osworld_2024; yang_synthesizing_2024). In sensitive domains such as healthcare and finance (tan_large_2024; jordan_machine_2015), privacy concerns further complicate data acquisition. Large generative models have emerged as capable synthetic data generators, producing annotated data for tasks like policy learning (xu_agenttrek_2024), Text2SQL (yang_synthesizing_2024), sentiment analysis (ye_zerogen_2022; li_synthetic_2023), and image classification (geng_unmet_2025). While synthetic data can improve real-world performance under scarcity, results vary widely depending on task and data quality (huang_survey_2025; geng_unmet_2025).
Can we distinguish between high-quality synthetic data that improves real-world task performance and low-quality data that offers little benefit, without any annotated real data and without costly model training? Crucially, increasing the size of synthetic datasets does not always lead to better downstream performance as it does with real data; in some cases, larger synthetic datasets can even degrade performance, exhibiting inverse scaling trends (geng_unmet_2025; li_synthetic_2023; setlur_rl_2024; gao_self-guided_2022; moller_parrot_2023). Therefore, selecting a synthetic dataset from a pool of datasets to train on to optimize downstream performance is important.
We introduce Synthetic Dataset Quality Estimation, or SynQuE, the problem of ranking multiple synthetic datasets by quality using only limited unannotated samples of real data. A synthetic dataset A is of higher quality than B if a model trained on A outperforms one trained on B on a real-world test set. This ability is crucial when real data annotation is costly or infeasible. For example, in text-to-SQL parsing, SynQuE helps select the synthetic dataset that yields better generalization from a small set of unannotated real queries. Similarly, for intelligent web agents, it identifies the synthetic interactions that produce agents performing best on real navigation tasks.
This work makes three main contributions. 1) We formalize the SynQuE problem and establish the first comprehensive benchmark. As part of this, we introduce and evaluate a suite of proxy metrics: computable scores that estimate a synthetic dataset’s quality using only a subset of unannotated real data. We adapt proxy metrics using established distributional measures such as mean distance to medoids (cox_directed_2021), diversity measures such as Proxy-A-Distance (ben-david_analysis_2006), and divergence measures such as MAUVE (pillutla_mauve_2021) — none have been systematically evaluated for the purpose of synthetic data selection. 2) We propose LLM-Evaluated Normalized Score (Lens), a novel proxy measure that leverages large language model (LLM) reasoning to create dataset rubrics that highlight difference between synthetic data and real data in language. 3) We conduct a comprehensive experiment across diverse domains in sentiment analysis, Text2SQL, web navigation, and image classification in order to evaluate how well these proxy metrics are able to select synthetic data to maximize performance on real test data.
Our empirical evaluation shows that SynQuE proxies exhibit moderate to strong correlation with real task performance across diverse domains including sentiment analysis, Text-to-SQL, image classification, and web navigation. While the best proxy varies by task, most can effectively predict downstream performance without any labeled real data, enabling practical synthetic data selection that outperforms indiscriminate synthetic data selection. We find that the reliability of proxies depends on task complexity, with higher variance on noisy data like synthetic images. Among the proxies, Lens, leveraging LLM reasoning and a principled debiasing strategy, consistently achieves superior gains on complex tasks like web navigation by capturing nuanced task details beyond embedding-based metrics. These results establish SynQuE as a robust framework for selecting high-quality synthetic data and motivate future work on stronger foundation model methods to characterize data and perform fine-grained, example-level data selection.
Figure 1: SynQuE uses synthetic data and unlabeled samples of real data to estimate synthetic data quality. Proxy scores are used to rank and select the datasets that lead to the best task performance
2 Related Work
Data synthesis
Synthesizing training data with generative models is a promising way to address data scarcity by leveraging instruction-following abilities (touvron_llama_2023; ouyang_training_2022) and vast pre-trained knowledge (long_llms-driven_2024; honovich_unnatural_2022; mishra_cross-task_2022). This has been applied across domains such as text classification  (ye_zerogen_2022), Text-to-SQL  (yang_synthesizing_2024; lei_spider_2024; li_codes_2024), planning  (sun_os-genesis_2024; hu_agentgen_2024; xu_agenttrek_2024; murty_nnetnav_2025), and computer vision  (geng_unmet_2025; li_bigdatasetgan_2022). While synthetic data often improves downstream models (liu_apigen_2024; ye_zerogen_2022), challenges remain in ensuring quality due to issues such as hallucination (huang_survey_2025), mode collapse (goodfellow_generative_2014; durall_combating_2020; shumailov_ai_2024), and counterfactual artifacts (li_synthetic_2023; yu_large_2023). Our work introduces the SynQuE framework for selecting synthetic data to maximize task performance without access to large amounts of real data and without training task models.
Evaluating data quality
Metrics like Proxy-$\mathcal{A}$-Distance ($\mathbf{PAD}$) estimate domain divergence by training classifiers to distinguish source and target data (ben-david_theory_2010; ben-david_analysis_2006; quinonero-candela_dataset_2022). While effective in some NLP tasks (elsahar_annotate_2019), $\mathbf{PAD}$ struggles in noisy, complex settings such as agent planning (he_webvoyager_2024). Diversity metrics like mean-distance-to-medoids ($\mathbf{MDM}$)(cox_directed_2021) and DCScore(zhu_measuring_2025) approximate the coverage of synthetic data. However, these measures do not capture common synthetic data factual inconsistencies from generative model hallucinations (huang_survey_2025; li_evaluating_2023; gunjal_detecting_2023), which can harm downstream performance (geng_unmet_2025; yu_large_2023; casco-rodriguez_self-consuming_2023; li_synthetic_2023; hataya_will_2023; briesch_large_2023). Other works focus on data selection using scaling laws by training a regression model on the results of short training runs to predict the best data mixture on a labeled validation set (liu_regmix_2025; magnusson_datadecide_2025). This approach, while effective, is computationally expensive and fundamentally requires labeled real data, which is unavailable in the SynQuE setting. Finally, another group of works try to distinguish between human and machine text using divergence measures (pillutla_mauve_2021) or LLM-as-a-judge (gu_survey_2025; krumdick_no_2025; zheng_judging_2023). Neither of these two techniques have been studied for synthetic data selection. Furthermore, the latter also suffers from limitations of generative models previously mentioned such as hallucination. Our work establishes the first systematic benchmark and results for synthetic data selection in a practical, fully unsupervised regime.
3 The Synthetic Dataset Quality Estimation Problem
We define the SynQuE problem and establish notation. Let $\mathcal{D}_{=}\{x_{^{(i)}},y_{^{(i)}}\}_{i=1}^{n}$ be a real dataset, which we assume is scarce and for which labels are unavailable for validation. Instead, we only have access to a small, unannotated collection of real-world inputs $\mathcal{U}_{=}\{x_{i}\}_{i=1}^{m\in\mathcal{D}}$. Our goal is to use $\mathcal{U}$ to estimate the quality of $K$ synthetic datasets $\{\mathcal{D}_{\mathrm{s}}^{(1)},\ldots,\mathcal{D}_{\mathrm{s}}^{(K)}\}$. These datasets might be generated by different methods and thus vary in quality. We aim to select the synthetic dataset $\mathcal{D}_{\mathrm{s}}^{*}$ that yields the best model performance on $\mathcal{D}$, without using labeled real data or training models on any synthetic dataset.
Consider a model $f(\cdot;\theta_{k})$ trained on the synthetic dataset $\mathcal{D}_{\mathrm{s}}^{(k)}$ with parameters $\theta_{k}$. Let $M(f,\mathcal{D}_{e})$ denote model $f$’s performance on the evaluation dataset $\mathcal{D}_{e}$, for instance task completion rate or accuracy. The ideal synthetic dataset $\mathcal{D}_{\mathrm{s}}^{*}$ satisfies:
$\mathcal{D}_{\mathrm{s}}^{*}=\mathrm{argmax}_{\mathcal{D}_{\mathrm{s}}^{(k)}}M\left(f\left(\cdot;\theta_{k}\right),\mathcal{D}_{e}\right)$ (1)
In practice, Eq 1 is infeasible because it requires labeled real data for evaluation and extensive model training across $K$ synthetic datasets. Instead, SynQuE seeks proxy metrics $Q(\mathcal{D}_{\mathrm{s}}^{(k)},\mathcal{U}_{)}$—computable using only the synthetic dataset $\mathcal{D}_{\mathrm{s}}^{(k)}$ and unannotated real samples $\mathcal{U}$—that correlate strongly with true downstream performance. The SynQuE problem thus reduces to designing or learning a proxy function $Q:(\mathcal{D}_{\mathrm{s}}^{(k)},\mathcal{U}_{)}\rightarrow\mathbb{R}$ such that a higher score indicates the synthetic dataset $\mathcal{D}_{\mathrm{s}}^{(k)}$ is more likely to produce better real-world task performance. This formulation enables efficient and effective use of synthetic data in low-resource or privacy-sensitive scenarios. For ease of exposition, we refer to the score produced by the proxy function $Q$ as the SynQuE score.
4 SynQuE Proxy Metrics
In this section, we introduce the suite of proxy metrics designed to solve the SynQuE problem. We study the potential of leveraging several distributional distance measures as a vehicle to quantify synthetic datasets quality. As the distributional distance measures have not been applied to synthetic dataset evaluation nor for tackling SynQuE, we adapt them to tackle SynQuE by embedding raw text or image data into representations, so distributional distances can be quantified through these traditional measures.
To address representation-based failure on evaluating long-horizon tasks, we introduce Lens  a LLM-based measure that operates directly on the raw data to be evaluated, therefore providing contextual understanding instead of sole depending on a universal embedding model. We introduce both our adapted representation-based metrics and LLM-based metrics in the following section.
4.1 Representation-Based Proxy Metrics
Mean Distance to Medoid
Our first proxy metric adapts Mean Distance to Medoid ($\mathbf{MDM}$), a measure of dataset diversity (cox_directed_2021; laliberte_distance-based_2010; lehman_abandoning_2011; risi_how_2009). The rationale is that high-diversity synthetic datasets may offer broader coverage in the embedding space and thus yield higher performance on real data. For SynQuE, $\mathbf{MDM}$ characterizes this diversity by measuring the sparsity of data points around medoids. Given $\mathcal{D}_{\mathrm{s}}$, we compute $N$ medoids using clustering algorithms such as kMedoids  [^1] https://pypi.org/project/kmedoids/. For each medoid $\tilde{x}^{M}$, we aggregate the Euclidean distances from all points within their corresponding cluster: $\mathbf{MDM}=\frac{1}{N}\sum_{i=1}^{N}d(x_{i}^{M},\tilde{x}^{M})$. Intuitively, if points within each cluster are centered around the medoid, $\mathbf{MDM}$ will be small, and therefore less diverse in the embedding space. In contrast, high-diversity synthetic datasets to have broader coverage, and therefore higher performance on real data. A higher $\mathbf{MDM}$ score suggests greater diversity, so we use it directly as the SynQuE score.
Maximum Mean Discrepancy
Next, we propose a proxy based on Maximum Mean Discrepancy ($\mathbf{MMD}^{2}$), a nonparametric test that assesses whether two samples originate from the same distribution (gretton_kernel_2012; borgwardt_integrating_2006; lu_improved_2022; li_generative_2015). As a SynQuE proxy, we use $\mathbf{MMD}^{2}$ to measure the discrepancy between the synthetic dataset and the real data distribution. Given $n_{\mathrm{s}}$ synthetic input samples $x_{i}\in\mathcal{D}_{\mathrm{s}}$ of size $n_{\mathrm{s}}$ and $m$ unannotated real input samples $y_{i}\in\mathcal{U}$, $\mathbf{MMD}^{2}$ quantifies the distance between these two empirical distributions in a reproducing kernel Hilbert space using a kernel function $k(\cdot)$.
$\mathbf{MMD}^{2}=\frac{1}{{m^{2}}\sum_{i,j=1}^{mk(x_{i},x_{j})+\frac{1}{{n_{\mathrm{s}}}^{2}}\sum_{i,j=1}^{n_{\mathrm{s}}}k(y_{i},y_{j})-\frac{2}{m_{n}{}_{\mathrm{s}}}\sum_{i,j=1}^{n_{\mathrm{s}},mk(x_{i},y_{j})}}}$ (2)
A smaller $\mathbf{MMD}^{2}$ score indicates the synthetic data distribution is closer to the real one. To maintain consistency with other proxies where higher is better, we use $\mathbf{MMD}^{2}$ as the SynQuE score.
Listing 1: Rubric prompt
You are shown samples
from datasets A and B.
Give up to 10 points
describing how dataset B
is similar to dataset A.
Samples from dataset A:
\${location A}
Samples from dataset B:
\${location B}
Listing 2: Scorer Prompt
Given similarities and differences between
datasets, how likely is the given sample
from dataset \${prediction}? Choose from very
unlikely, unlikely, unsure, likely, and very
likely.
Similarities: \${similarities}
Differences: \${differences}
Sample: \${sample}
Table 1: Simplified Lens prompt templates. Appendix C contains detailed prompts for each task.
Proxy-A-Distance
Our third representation-based proxy adapts Proxy-A-Distance ($\mathbf{PAD}$), a discriminative measure from domain adaptation that quantifies the divergence between two distributions (ben-david_analysis_2006; elsahar_annotate_2019). The $\mathbf{PAD}$ proxy measures how well a classifier can discriminate between samples from the synthetic dataset $\mathcal{D}_{\mathrm{s}}$ and real dataset $\mathcal{D}$. We compute $\mathbf{PAD}=1-2\mathcal{E}(G)$ by training a binary domain classifier $G:x\rightarrow\left[0,1\right]$ (e.g., a linear SVM, a multi-layer perceptron) to distinguish between synthetic and real inputs (i.e. we do not assume access to labels), where the error of the classifier $\mathcal{E}$ can be computed as:
$\mathcal{E}\left(G\right)=1-\frac{1}{n_{\mathrm{s}}+m}\sum_{x_{i}\in\mathcal{D}_{\mathrm{s}},\mathcal{U}\left|G\left(x_{i}\right)-\mathbb{I}\left(x_{i}\in\mathcal{D}_{\mathrm{s}}\right)\right|}$ (3)
A higher classification error implies the datasets are not easily separable, indicating lower divergence and thus higher synthetic data quality. For consistency with other divergence metrics, use $\mathbf{PAD}$ as the SynQuE score.
Mauve
Our final representation-based proxy is an adaptation of MAUVE (pillutla_mauve_2021), a metric that quantifies the divergence between text distributions. Mauve summarizes both Type I and Type II errors. Given a synthetic data distribution Q and a real data distribution P, Type I error means Q generates text that is unlikely under P (unrealistic samples), while Type II error means Q fails to generate text that is plausible under P (lacks diversity). Mauve captures both error types in a single score, which approaches 1 as the distributions become more similar. Since a higher score indicates better alignment with real data, we use the MAUVE score directly as the SynQuE score.
4.2 LLM-Evaluated Normalized Score (Lens)
The representation-based proxies described so far rely on high-quality continuous representations of inputs. In low-resource settings where such representations are unavailable, or in long-horizon settings where it is intractable to compress a long sequence of observations and states into a compact, fixed-size representation, these representation-based proxies may fall short. To address this, we introduce LLM-Evaluated Normalized Score (Lens), a novel method that leverages LLMs as zero-shot discriminators. Lens first derives a language rubric describing the similarities and differences between samples of unannotated real data $\mathcal{U}$ and inputs from the synthetic dataset $\mathcal{D}_{\mathrm{s}}$. A subsequent (potentially smaller) LLM then scores how likely each synthetic example is to belong to the real dataset, guided by the rubric. The average score across the synthetic dataset is used as the final SynQuE score. The intuition is similar to that behind $\mathbf{PAD}$: a higher classification error by the rubric-guided scorer implies higher synthetic data quality. We now detail how Lens is computed.
Rubric compilation
Given real input samples $\mathcal{U}$, we collect an equal number of samples $\mathcal{U}_{\mathrm{s}}$ from the synthetic dataset $\mathcal{D}_{\mathrm{s}}$. Both collections are given to a reasoning LLM (e.g. DeepSeek R1 or o4-mini) to generate three sets of characteristic descriptions: commonalities ($C$), differences of real from synthetic ($C_{,\mathrm{s}}$), and differences of synthetic from real ($C_{\mathrm{s},}$). Listing LABEL:lst:prompt_rubric shows a simplified rubric compilation prompt template. Our design is backed by the principled idea of approximating domain divergence through discriminator error (ben-david_analysis_2006): Lens’s scoring is motivated by $\mathbf{PAD}$, where the error of a classifier (here, the LLM-based scorer) reflects the distance between distributions. Unlike $\mathbf{PAD}$, however, Lens does not require a pretrained encoder to map samples into fixed-length representations; instead, it operates directly on the native data format and characterizes differences using language rubrics.
Principled Debiasing and Scoring
We now describe how to compute the score of a synthetic dataset. A key challenge of scoring is in mitigating LLM biases. We identified three primary sources:
  1. 1.
    Order Bias: The set of differences an LLM derives when comparing A to B can differ significantly from when comparing B to A.
  2. 2.
    Label Bias: When asked how likely an example $x$ belongs to A or B, an LLM may score both as “very likely”, a contradiction.
  3. 3.
    Score Bias: LLMs may have an inherent preference for certain score values (e.g., “likely”) regardless of the input.
To address these systematically, we employ a minimal design involving four scoring permutations for each sample. We denote the LLM scoring function as $g_{\mathcal{D}\mid C}$, which outputs a score (0-4) for how likely an example $x$ belongs to dataset $\mathcal{D}$ given characteristic descriptions $C$.
First, to mitigate score bias, we compute baseline scores by averaging the LLM’s judgments on real inputs $x\in\mathcal{U}$ for each of the four permutations. For instance, the baseline for scoring an example as real, given the description of how synthetic differs from real, is:
$z_{\mid C_{\mathrm{s},}}=\mathbb{E}\left[{g_{\mid C_{\mathrm{s},}}(x)}\right]\approx\frac{1}{n}\sum_{i=1}^{ng_{\mid C_{\mathrm{s},}}(x_{i})}$ (4)
We then compute a score-debiased score $h$ for each synthetic sample by normalizing its raw score against this baseline:
$h_{\mid C_{\mathrm{s},}}=\frac{g_{\mid C_{\mathrm{s},}}(x)}{\mathrm{max}(\epsilon,z_{\mid C_{\mathrm{s},}})}$ (5)
Here, $\epsilon$ is a small constant to avoid division by zero. Intuitively, this scores-debiased score expresses how much the LLM scores the example compared to how it usually scores real examples. Next, to compute a label-debiased score, we normalize the LLM’s preference for the “real” label over the “synthetic” label for each synthetic example:
$p_{\mid C_{\mathrm{s},}}=\frac{h_{\mid C_{\mathrm{s},}}}{h_{\mid C_{\mathrm{s},}}+h_{\mathrm{s}\mid C_{\mathrm{s},}}+\epsilon}$ (6)
Finally, to create an order-debiased score, we average the label-debiased scores obtained using both sets of difference descriptions ($C_{\mathrm{s},}$ and $C_{,\mathrm{s}}$):
$\hat{p}(x)=\frac{1}{2}\left[p_{\mid C_{\mathrm{s},}}(x)+p_{\mid C_{,\mathrm{s}}}(x)\right]$ (7)
The final Lens score for a synthetic dataset is the empirical mean of these fully debiased scores across all its samples:
$\mathrm{\textsc{Lens}}(\mathcal{D}_{\mathrm{s}})=\mathbb{E}\left[{\hat{p}(x)}\right]=\frac{1}{n}\sum_{i=1}^{n_{\mathrm{s}}}\hat{p}(x_{i})$ (8)
5 Experiments
We choose four diverse tasks spanning different machine learning domains to examine how well each candidate proxy metric extrapolate to real data performance on tasks with varying complexities and modalities. For Lens, we incorporate Deepseek-R1 to generate $10$ points about similar and different characteristics $C_{\mathrm{s},}$ between synthetic data samples $\mathcal{U}_{\mathrm{s}}$ and real data samples $\mathcal{U}$. We then use Qwen2.5-32B-Instruct (qwen_qwen25_2025) and 8B to score synthetic examples according to the rubric. For image domain, we use OpenAI o4-mini to compile rubrics and Qwen2.5-VL-32B-Instruct to score. For representation-based metrics $\mathbf{PAD},\mathbf{MDM},\text{and }\mathbf{MMD}^{2}$, we use state-of-the-art qte-Qwen2-7B-Instruct to embed text inputs and E5-V (jiang_e5-v_2024) to embed image inputs for proxy scoring. We use XGBoost  (chen_xgboost_2016) to compute $\mathbf{PAD}$ and polynomial kernel for $\mathbf{MMD}^{2}$  (gretton_kernel_2012). We include additional kernel ablations in Appendix 10. We use the official release [^2] https://pypi.org/project/mauve-text/ from Mauve, with the default hyperparameter setting for Mauve calculation.
We also include an experiment with perplexity-based metric as additional baseline in Text2SQL (see  table 5). Perplexity, inspired by scaling law methods (magnusson_datadecide_2025; liu_regmix_2025), fine-tunes a model on each synthetic data set and measures its perplexity on the unannotated real data subset; a lower perplexity is expected to indicate higher quality.
We use Pearson (pearson_vii_1997) and Spearman rank (spearman_proof_1904) correlation coefficients to measure how strongly task performance and proxy scores are related. Pearson focuses on predictability by capturing linear relationships, while Spearman focuses on trend by evaluating whether the relationship between variables is consistently increasing or decreasing (i.e., monotonic), regardless of the exact shape. To reduce variance in correlation analysis across different sample subsets, for all tasks, we construct subsets $\mathcal{U}$ by sampling with five different seeds. Final correlation scores are averaged across seeds.
Table 2: Top-3 task performance for all candidate proxies of SynQuE. Top-3 task performance is computed by averaging task performance of synthetic datasets chosen using each proxy metric. Improvements are calculated based on increase over average performance of all synthetic datasets.
Top-3 Ranked Average Task Performance across Proxy Metrics
Tasks Test Lens 7B Lens 32B $\mathbf{PAD}$ $\mathbf{MMD}^{2}$ $\mathbf{MDM}$ Mauve
mean debiased biased debiased biased
Sentiment 49.6 50.5 +0.8 51.2 +1.6 52.0 +2.4 51.0 +1.4 55.3 +5.7 54.7 +5.1 54.2 +4.6 54.6 +4.9
Text2SQL
Computer 45.8 46.6 +0.7 46.4 +0.6 48.3 +2.5 46.3 +0.5 48.3 +2.5 47.4 +1.6 48.2 +2.3 48.3 +2.5
Apps 30.4 34.7 +4.4 36.3 +5.9 33.8 +3.4 35.2 +4.9 33.5 +3.2 38.4 +8.1 33.9 +3.5 38.4 +8.1
Movies 37.3 41.0 +3.7 41.4 +4.1 43.8 +6.5 44.7 +7.4 44.2 +6.9 43.0 +5.7 46.9 +9.6 44.6 +7.3
Average 37.8 40.8 +2.9 41.4 +3.5 42.0 +4.1 42.1 +4.3 42.0 +4.2 42.9 +5.1 43.0 +5.2 43.8 +6.0
Image
Split 1 57.2 57.3 +0.2 56.7 -0.4 56.4 -0.8 53.4 -3.7 55.9 -1.3 57.3 +0.1 57.0 -0.1 56.0 -1.1
Split 2 55.8 55.3 -0.4 55.8 +0.0 56.2 +0.4 56.0 +0.2 55.4 -0.4 54.5 -1.3 54.8 -1.0 56.3 +0.5
Split 3 57.7 59.1 +1.4 58.7 +1.0 60.2 +2.5 57.0 -0.7 58.2 +0.6 64.1 +6.5 52.2 -5.5 58.4 +0.7
Average 56.9 57.2 +0.4 57.0 +0.2 57.6 +0.7 55.5 -1.4 56.5 -0.4 58.6 +1.8 54.7 -2.2 56.9 +0
WebNav 25.8 26.5 +0.7 26.3 +0.5 26.3 +0.5 26.0 +0.2 25.7 -0.1 26.5 +0.7 25.8 -0.1 26.3 +0.5
Sentiment Analysis
Recent work shows LLMs overfit widely-used datasets due to data contamination (balloccu_leak_2024; sainz_nlp_2023; oren_proving_2023). To mitigate this, we evaluate on a domain-specific financial tweets sentiment dataset [^3] https://huggingface.co/datasets/zeroshot/twitter-financial-news-sentiment. We create 32 synthetic class-balanced datasets (998 samples each) using eight prompt types: zero-shot, zero-shot with background knowledge, with train-time or test-time stock ticker info, and few-shot variants. We use Qwen2.5-7B-Instruct, Qwen2.5-32B-Instruct, Llama3.1-8B-Instruct, and Llama3.3-70B-Instruct (grattafiori_llama_2024) for each prompt type. Background knowledge uses detailed guideline instructions for better task alignment. Stock tickers are sampled one at a time for synthesis. Details are in appendix C. We train task models using XGBoost and evaluate F1 score on a 2,388-item test set. Rubrics are compiled by randomly sampling 200 points from each real and synthetic dataset.
Text2SQL
We evaluate SynQuE on Text2SQL using three DBs from the BIRD benchmark (Movies, App Store, Computer Students — we the last two as Apps and Computers) (li_can_2023). We synthesize 1,000 data points with 4 prompt types: zero-shot with background knowledge (guidelines and schema), zero-shot with test-time info (random table rows), and few-shot (three examples). Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct models are used for dataset generation. Task models are finetuned from Qwen2.5-Coder-1.5B-Instruct following CodeS [^4] https://github.com/RUCKBReasoning/codes and evaluated using execution accuracy on the real test set. For rubrics, we sample 30 points per synthetic and real dataset. Real data sizes are 60, 69, and 164 for Apps, Computers, and Movies respectively.
Table 3: Spearman (left) and Pearson (right) correlation scores of SynQuE proxy metrics. Lens uses a fraction of samples for rubric compilation except for Web Navigation tasks.
Tasks Lens 7B Lens 32B $\mathbf{PAD}$ $\mathbf{MMD}^{2}$ $\mathbf{MDM}$ Mauve
debiased biased debiased biased
Sentiment .25 .33 .26 .17 .38 .26 .24 .23 .53 .65 .45 .67 .68 .85 .53 .57
Text2SQL
Computer .19 .13 .10 .10 .41 .45 .18 .23 .46 .69 .33 .85 .39 .63 .24 .78
Apps .38 .37 .42 .49 .46 .40 .55 .61 .43 .42 .53 .79 .44 .56 .74 .52
Movies .41 .50 .41 .26 .50 .46 .56 .47 .50 .64 .38 .46 .61 .41 .65 .68
Average .33 .33 .31 .28 .46 .43 .43 .44 .46 .58 .41 .70 .48 .53 .55 .66
Image
Split 1 -.18 -.19 -.30 -.35 -.28 -.28 -.68 -.67 -.06 -.05 .66 .52 -.37 -.27 -.04 -.20
Split 2 .02 -.04 .14 .05 .20 .05 .20 .05 .20 .31 .09 .17 .03 -.32 .03 .13
Split 3 -.10 -.01 -.15 -.03 .31 .33 .37 .44 .02 -.15 .26 .21 -.54 -.76 .46 .34
Average -.09 -.08 -.10 -.11 .08 .03 -.04 -.06 .05 .04 .33 .30 -.30 -.45 .15 .09
WebNav .15 .17 .11 .18 .15 .15 .08 .09 .11 .08 -.02 .06 -.11 -.08 -.09 -.10
Table 4: Division of the 15 selected ImageNet classes into three 5-class splits for image classification tasks. Each row corresponds to one split used in our experiments.
Splits ImageNet classes
1 bra, mask, lion, cloak, tank
2 hammer, backpack, stage, throne, tray
3 plate, desk, kimono, shield, church
Image Classification
In addition to text-only settings, we evaluate SynQuE on image classification using synthetic datasets curated from unmet-promise [^5] https://huggingface.co/datasets/scottgeng00/unmet-promise. These datasets are created with different prompts using Stable Diffusion 1.1 and 1.5: label, label plus physical relation, and label plus background description (geng_unmet_2025). Images are mapped to ImageNet classes (deng_imagenet_2009) via caption analysis C, then filtered with a vision-language model to remove noisy labels. Data cleaning details are provided in Appendix appendix C. The final set includes 15 classes with 300 images each. Due to limited samples, the 15-class task is split into three 5-class tasks (table 4). We train ResNet-50 (he_deep_2015) from scratch for 50 epochs, early stopping on 10% validation data. Evaluation uses mean reciprocal rank (MRR) for finer performance measurement. Rubrics are constructed from 100 sampled images per real and synthetic data, consistent across SynQuE methods.
Web Navigation
Our fourth task evaluates SynQuE on agentic web navigation planning using WebVoyager (he_webvoyager_2024) and synthetic data from NNetNav (murty_nnetnav_2025). Inputs include task objectives, current step observations (accessibility tree), and past actions; targets are actions leading to task success. WebVoyager has 15 websites; we exclude Google Flights and Booking which are no longer feasible (zhou_proposer-agent-evaluatorpae_2024; murty_nnetnav_2025), leaving 13 sites with 557 tasks. Each site forms a test domain split into 5 synthetic subsets. Models are fine-tuned with LoRA (hu_lora_2021) on Qwen2.5-7B-Instruct. We use all synthetic and 20 real samples per method.
5.1 Results Analysis
Table 5: Spearman (left) and Pearson (right) correlations with Perplexity scoring on the BIRD Text2SQL benchmark
Perplexity
Computers -.31 -.33
Apps -.25 -.29
Movies .24 .31
Average -.11 -.10
SynQuE proxies correlate with task performance and improve selection.
table 3 shows that SynQuE proxy metrics demonstrate moderate to strong correlation with downstream task performance. To show the practical utility of these proxies, we simulate selecting the top 3 datasets based on each metric’s score and compare their average task performance against the mean performance of all available synthetic datasets. As shown in table 2, nearly all proxy metrics significantly improve dataset selection over selecting synthetic datasets non-discriminately (i.e. uniform selection). This demonstrates that SynQuE is an effective framework for maximizing real-data performance, despite not having access to labeled real data and only a limited sample of real data. We also conduct an experiment with Perplexity on Text2SQL, to examine the effectiveness as a potential proxy. As shown in table 5, it correlates poorly with even text data, therefore we conclude that scaling methods would not work under our setting, where no annotated data is available for evaluation to build a regression model that predicts the best data mixture.
Performance on ambiguous image data shows high variance.
The synthetic image classification data contains significant visual variability and label ambiguity, especially in Split 1 between classes like “stage” and “throne” (fig. 2(a), fig. 2(b)). This confuses most proxy metrics, resulting in inconsistent correlations across splits (table 3). However, table 2 shows that when used for selection, several proxies (e.g. debiased Lens 32B, $\mathbf{MMD}^{2}$) still improve average task performance.
Table 6: Spearman (left) and Pearson (right) correlations with different number of real samples for scoring Text2SQL. Lens uses debiased 32B scoring. Note that results in table 3 use 30 real samples.
$\left|\mathcal{U}\right|=25$ $\left|\mathcal{U}\right|=50$
Lens PAD $\mathbf{MMD}^{2}$ $\mathbf{MDM}$ Mauve Lens $\mathbf{PAD}$ $\mathbf{MMD}^{2}$ $\mathbf{MDM}$ Mauve
Computers .22 .22 .36 .65 .28 .80 -.39 -.65 .19 .65 .64 .38 .51 .78 .34 .87 -.39 -.62 .87 .78
Apps .29 .48 -.20 -.23 .40 .77 -.44 -.56 .65 .70 .64 .66 .68 .76 .57 .80 -.44 -.56 .32 .65
Movies .33 .48 .33 .36 .52 .57 -.60 -.40 -.14 -.35 .43 .49 .57 .37 .81 .56 -.67 -.45 .81 .60
Average .28 .39 .16 .26 .40 .71 -.47 -.54 .23 .33 .57 .51 .59 .64 .57 .74 -.50 -.55 .67 .68
Using more real samples improves correlation.
As shown in table 6, increasing the number of unannotated real samples $m$ consistently leads to stronger correlations for all proxy metrics. This indicates that even a modest increase in available real-world data can significantly improve the reliability of synthetic data quality estimation.
Lens excels on complex, long-horizon tasks.
As shown in both tables, the 32B debiased Lens is the only proxy that consistently achieves positive correlation and improves top-3 task performance across all tasks and splits. Its advantage is particularly pronounced in web navigation, a complex planning task where representation-based metrics struggle. Lens leverages LLM reasoning over rich, structured inputs like accessibility trees to generate interpretable rubrics. For instance, an example characteristic point for the website “Wolfram Alpha” is: “Dataset B tasks focus on data retrieval (e.g. temperature anomalies, moon phases) while Dataset A emphasizes applied computational problem-solving”. These specific nuances in long text (e.g. instructions, state observations) are difficult to capture using general-purpose dense vector embedders, which explains why Lens outperforms representation-based proxy metrics on complex, abstract tasks like web navigation. This method does exhibit weaker correlation in image classification. We hypothesize that this stems from the inability of VLMs to capture meaning characteristics descriptions in batches of images during rubric generation, and VLM rubric generation will likely improve as VLMs improve in quality.
5.2 Ablation Studies and Cost Analysis
Cost-Effectiveness of SynQuE Proxies
A key motivation for SynQuE is efficiency. Our representation-based proxies require a one-time embedding computation, after which scoring all datasets is nearly instantaneous (e.g., 19 seconds for MMD on 32 datasets). Lens, using modern LLM serving frameworks, is also highly efficient, taking $\sim 15$ seconds per dataset with a 32B model on a H200 GPU. In contrast, perplexity-based data selection, inspired by scaling-law studies (liu_regmix_2025), require training many (e.g., 512 1M models used in their experiment) small models on the mixture of all synthetic datasets, a significantly more costly procedure, yet yield weaker correlations (table 3). This highlights the practical advantage of the SynQuE framework.
Larger scorers lead to stronger correlations
We find that larger scoring models yield stronger correlations between Lens and task performance, as shown in table 3. Intuitively, this is expected because larger models generally possess more robust instruction-following and reasoning capabilities, enabling them to better assess data quality and align proxy scores with downstream performance.
Lens is robust to preferential bias in LLM training data.
To address concerns that an LLM evaluator might favor data it generated, we tested Lens with different scoring models on data generated by Qwen2.5. The results, detailed in Appendix table 7, show that performance is consistent across evaluators, including those distinct from models used to generate the synthetic data, indicating minimal preferential bias.
Principled debiasing and rubrics are critical for Lens.
As illustrated in  table 3, the correlations between Lens and task performance consistently increase when debiasing is applied, indicating that raw scores may be systematically biased and do not reliably reflect true data quality. Once debiasing is introduced, the correlation becomes strongly positive and consistently outperforms the biased scores. This demonstrates that debiasing effectively corrects for these systematic errors and aligns Lens scores with actual task performance. Further ablations show that using a rubric consistently improves correlation over a zero-shot baseline (Appendix table 8), and that 10 rubric points generally offers the best trade-off between specificity and generality (Appendix table 9).
6 Conclusions, Limitations and Future Work
We formalized the SynQuE problem of ranking synthetic datasets by their impact on real-world task performance using limited unannotated real data. Our comprehensive evaluation established that various proxies can reliably predict downstream performance, offering a cost-effective alternative to full model training. We proposed Lens, a novel proxy leveraging LLM reasoning and principled debiasing, which consistently outperforms others on challenging, long-horizon tasks. Overall, SynQuE offers a robust framework for synthetic data selection when labeled real data is scarce.
SynQuE assume that real data is scarce, a setting not all deployments face. While Lens performs well on the complex tasks studied, its effectiveness should be validated on more diverse tasks. Additionally, we experiment with limited-size LLMs due to resource constraints. Future work should explore 1) scaling Lens to larger sizes and different architectures, especially strong VLMs, to assess generality and improvements. 2) using rubric feedback to guide LLMs in synthesizing more realistic data, and 3) developing fine-grained, example-level proxy use to directly improve task model training.
Appendix A Visualization of ambiguous image
(a) Sample synthetic images of class ”stage”
(b) Sample synthetic images of class ”throne”
Figure 2: Visualization of synthetic images from second split for classes (a) ”stage” and (b) ”throne”
Appendix B Additional Ablations
B.1 Ablation study on Lens preferential bias
Table 7: We use two non-Qwen scoring models on tasks where synthetic data was generated by Qwen2.5 models. The consistent positive correlations of the debiased scores demonstrate that Lens is robust to this potential bias.
LENS (Granite-8B) Debiased Biased
Task Spearman Pearson Spearman Pearson
Sentiment .21 .30 -.04 -.02
Text2SQL .22 .20 .19 .23
LENS (Ministral-8B) Debiased Biased
Task Spearman Pearson Spearman Pearson
Sentiment .28 .39 .52 .31
Text2SQL .32 .29 .16 .16
B.2 Ablation study on Lens rubric design
Table 8: LENS with vs. without a rubric on sentiment analysis, showing the rubric’s effectiveness. We use Qwen2.5-7B-Instruct for scoring.
LENS Model w/ Rubric Spearman Pearson
Qwen2.5-7B-Instruct NO 0.23 0.21
Qwen2.5-32B-Instruct NO 0.32 0.13
Qwen2.5-7B-Instruct YES 0.25 0.33
Qwen2.5-32B-Instruct YES 0.38 0.26
Table 9: Bottom: Varying the number of rubric points, where 10 points provides a good balance.
# Rubric Points Sentiment Analysis Text2SQL (Avg)
(Debiased) Spearman Pearson Spearman Pearson
5 0.14 0.19 0.38 0.30
10 0.25 0.33 0.33 0.33
15 -0.06 0.05 0.23 0.21
B.3 Ablation study on MMD2 kernel functions
Table 10: We report Pearson / Spearman correlation coefficients. The Laplacian kernel performs best on text tasks, but all kernels show limited effectiveness on more complex domains.
Task Polynomial RBF Laplacian Linear Sigmoid
Sentiment .67 / .45 .67 / .45 .78 / .60 .67 / .45 .67 / .46
Text2SQL (avg) .51 / .43 .52 / .44 .70 / .51 .52 / .44 .53 / .44
Image (avg) .22 / .30 .22 / .30 .21 / .30 .22 / .30 .22 / .30
WebNav .06 / -.02 .06 / -.01 .05 / -.04 .06 / -.02 .05 / -.03
Appendix C Additional Experimental Details
C.1 Data Curation
Sentiment analysis
For sentiment analysis, we used a cleaned version of the original validation split as the test set, removing URLs from the data. Synthetic samples are generated and validated to ensure each output is non-empty, alphabetic ’headline’ and a ’sentiment’ label restricted to the values ’0’, ’1’, or ’2’. Only samples meeting these structural and content criteria were retained for downstream analysis.
Text2SQL
During Text2SQL data synthesis, we validate generated question–SQL pairs by ensuring both fields are non-empty, contain alphanumeric content, and are formatted as a dictionary with ’question’ and ’SQL’ keys. For non-zero-shot generations, SQL queries are executed against the target database; pairs failing execution are discarded. This process enforces correct structure, meaningful content, and SQL executability.
Image classification
We classify each image in the unmet-promise dataset into one of three categories: label_only, label_relation, or label_background, based on its caption. Captions are lowercased and stemmed. If a class-specific background keyword appears in the caption, the image is assigned to label_background. If a relation keyword is present, it is assigned to label_relation. Images not matching either are assigned to label_only. We balance the number of samples per category and class, and store the processed data for downstream tasks. To further ensure fidelity of images, we use Qwen2.5-VL-7B-Instruct to filter noisy images. Prompt used for filtering is provided in listing  LABEL:lst:filter_prompt.
Web navigation
We first group individual trajectories into tasks. Then we use different seeds to create 5 disjoint subsets with equal amount of tasks. We fine-tune using LoRA rank of 64 and the open-instruct  [^6] https://github.com/allenai/open-instruct fine-tuning codebase.
C.2 SynQuE Scoring
For $\mathbf{PAD}$, we reserve 20% of all embeddings as a holdout test set to train the classifier, and compute the classification error on this set to obtain the final $\mathbf{PAD}$ score. For $\mathbf{MMD}^{2}$, we generate the proxy score using a polynomial kernel  [^7] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.pairwise.polynomial_kernel.html with degree 3 and Coef0 parameter 1. For $\mathbf{MDM}$, we use Fasterpam (schubert_faster_2019) to compute $k$ medoids in the embeddings, setting $k$ to 3 for sentiment analysis and 5 for other tasks, using Euclidean distance for clustering; $\mathbf{MDM}$ is then calculated by averaging the Euclidean distance from each data point to its corresponding medoid. For Lens, the prompts used for rubric compilation and scoring are provided in section C.5.3 and section C.5.4, respectively. For Mauve calculation, we use the default hyperparameter, the same embedding model as the other representation-based proxies (qte-Qwen2-7B-Instruct), and the scaling factor is 5. For Perplexity calculation in Text2SQL, we first reformat each Text2SQL question as a string: "question": <sample>. We then use Qwen2.5-7B to compute the perplexity, considering only the tokens corresponding to the question text (i.e., the ¡sample¿ portion), and excluding the prompt tokens (such as "question":).
C.3 Standard deviation of correlation coefficients
Table 11: Standard deviation of Spearman (left) and Pearson (right) correlation coefficients of SynQuE proxy metrics across 5 seeds. $\mathbf{MDM}$ only scores on synthetic datasets therefore no change in input data.
Tasks Lens 7B Lens 32B PAD MMD MDM
debiased biased debiased biased
Sentiment .23 .17 .12 .14 .09 .11 .07 .07 .02 .02 .02 .02 .00 .00
Text2SQL
Computer .48 .39 .38 .44 .21 .21 .18 .14 .04 .08 .03 .08 .00 .00
Apps .38 .34 .25 .21 .22 .22 .22 .14 .36 .48 .05 .15 .00 .00
Movies .32 .32 .33 .24 .22 .24 .11 .14 .07 .13 .05 .14 .00 .00
Average .39 .35 .32 .30 .22 .22 .17 .14 .15 .23 .04 .12 .00 .00
Image
Split 1 .47 .48 .51 .51 .39 .38 .28 .22 .08 .06 .00 .02 .00 .00
Split 2 .43 .47 .43 .42 .40 .41 .43 .53 .11 .08 .00 .04 .00 .00
Split 3 .52 .39 .47 .39 .40 .42 .29 .22 .07 .07 .00 .05 .00 .00
Average .47 .44 .47 .44 .40 .40 .33 .32 .09 .07 .00 .04 .00 .00
WebNav .15 .17 .11 .13 .09 .11 .15 .19 .04 .02 .02 .02 .00 .00
C.4 Example Rubrics
Listing 3: Example characteristic descriptions $C_{\mathrm{s},}$
"Dataset B consistently specifies the analyst behind actions...",
"Dataset B maintains strict financial focus without political or
entertainment tangents present in A...",
"Dataset B entries always directly connect stock movements to specific
analyst actions...",
"Dataset B shows more frequent price target amount disclosures...",
"Dataset B uses standardized financial terminology consistently...",
"Dataset B maintains neutral tone in earnings reports...",
"Dataset B focuses exclusively on institutional analyst
perspectives...",
"Dataset B headlines strictly follow ’[Analyst] [Action] on [Ticker]
[Rationale]’ structure...",
"Dataset B contains no social media tags/hashtags...",
"Dataset B shows higher frequency of ETF coverage..."
Listing 4: Example characteristic descriptions $C_{,\mathrm{s}}$
"Dataset B includes headlines without stock tickers ...",
"Dataset B contains non-financial news ...",
"Dataset B incorporates social media-style commentary...",
"Dataset B includes international/non-English company names...",
"Dataset B references non-institutional analysts/sources...",
"Dataset B features headlines about dividends...",
"Dataset B includes legal/regulatory actions unrelated to markets...",
"Dataset B uses technical trading jargon...",
"Dataset B contains macroeconomic commentary without stock links...",
"Dataset B includes non-company-specific index/currency forecasts..."
C.5 LLM Usage
Model serving
We use vLLM (kwon_efficient_2023) to serve open source models such as for data synthesis and dataset scoring. For Llama3.3-70B-Instruct model, we use Ollama  [^8] https://ollama.com/ Q4_K quantized version to construct synthetic datasets for sentiment analysis. We use 2 * Nvidia A40 48GB GPUs for other models for synthesis and scoring. To improve LLMs’ generation throughput, we use vLLM’s batched inference feature and enable prefix-caching to further improve generation efficiency.
LLM hyperparameter
For Lens rubric compilation and scoring, we set temperature to 0 and top_p to 0.95.
C.5.1 Data Synthesis Prompts
Listing 5: Zero-shot prompt used for sentiment analysis dataset generation.
Generate three realistic financial news headlines for sentiment analysis.
Guidelines for Generating Headlines:
Sentiment Labeling:
Each headline must be assigned a sentiment label based on its tone:
- Bearish (0): Indicates negative sentiment about a stock or market trend.
- Bullish (1): Indicates positive sentiment about a stock or market trend.
- Neutral (2): Indicates neutral or informational tone.
Now, generate three new financial news headlines following these guidelines. Please use JSON format and generate one type of each sentiment label (0, 1, 2) in your response.
Listing 6: Zero-shot with background knowledge prompt used for sentiment analysis dataset generation.
Generate three realistic financial news headlines about stock tickers following real-world financial reporting for sentiment analysis.
Guidelines for Generating Headlines:
1. Format & Style:
- Headlines must be concise and mimic real financial news.
- Use sentence case formatting (capitalize only the first word and proper nouns).
- Some headlines should start with a stock ticker (e.g., \$AAPL -), while others should begin with the company name or a broader market trend.
2. Ticker Inclusion:
- At least one headline should include a stock ticker (e.g., \$TSLA - or \$NVDA -).
- Some headlines should refer to companies by name instead of tickers (e.g., "Alphabet and Meta see price targets cut at Barclays").
3. Common Financial Themes:
Ensure headlines reflect realistic financial news topics,
including:
- Stock downgrades/upgrades
- Price target adjustments
- Market trends/economic outlook
- Company performance concerns
- Company news
- Company announcements
- Company events
4. Source Attribution:
- When relevant, mention an investment firm, analyst, or research group (e.g., Morgan Stanley, Barclays, Oppenheimer).
- Do not fabricate research firms-use only well-known institutions.
5. Sentiment Labeling:
Each headline must be assigned a sentiment label based on its tone:
- Bearish (0): Indicates negative sentiment about a stock or market trend.
- Bullish (1): Indicates positive sentiment about a stock or market trend.
- Neutral (2): Indicates neutral or informational tone.
Sentiment Labeling:
Each headline must be assigned a sentiment label based on its tone:
- Bearish (0): Indicates negative sentiment about a stock or market trend.
- Bullish (1): Indicates positive sentiment about a stock or market trend.
- Neutral (2): Indicates neutral or informational tone.
Now, generate three new financial news headlines following these guidelines. Please use JSON format and generate one type of each sentiment label (0, 1, 2) in your response.
Listing 7: Zero-shot with background and stock ticker information prompt used for sentiment analysis dataset generation.
Generate three realistic financial news headlines about stock tickers following real-world financial reporting for sentiment analysis.
Guidelines for Generating Headlines:
1. Format & Style:
- Headlines must be concise and mimic real financial news.
- Use sentence case formatting (capitalize only the first word and proper nouns).
- Some headlines should start with a stock ticker (e.g., \$AAPL -), while others should begin with the company name or a broader market trend.
2. Ticker Inclusion:
- At least one headline should include a stock ticker (e.g., \$TSLA - or \$NVDA -).
- Some headlines should refer to companies by name instead of tickers (e.g., "Alphabet and Meta see price targets cut at Barclays").
3. Common Financial Themes:
Ensure headlines reflect realistic financial news topics, including:
- Stock downgrades/upgrades
- Price target adjustments
- Market trends/economic outlook
- Company performance concerns
- Company news
- Company announcements
- Company events
4. Source Attribution:
- When relevant, mention an investment firm, analyst, or research group (e.g., Morgan Stanley, Barclays, Oppenheimer).
- Do not fabricate research firms-use only well-known institutions.
5. Sentiment Labeling:
Each headline must be assigned a sentiment label based on its tone:
- Bearish (0): Indicates negative sentiment about a stock or market trend.
- Bullish (1): Indicates positive sentiment about a stock or market trend.
- Neutral (2): Indicates neutral or informational tone.
Now, generate three new financial news headlines about stock tickers: {stock_ticker} following these guidelines. Please use JSON format and generate one type of each sentiment label (0, 1, 2) for diversity.
C.5.2 Data Cleaning Prompts
Listing 8: Prompt used for filtering noisy synthetic images
You are a helpful assistant that filters out an image. You will be given an image and its corresponding text caption.
You should return true if the primary object in the image is not a \${label} in common sense. Return false otherwise.
Image:
{image}
Caption:
{caption}
C.5.3 Lens Rubric Compilation Prompts
Listing 9: Rubric compilation prompt used in sentiment analysis (commonalities)
You are a world class data analyst on financial news headline. You will be given some financial news headline samples from dataset A and dataset B. Based on the provided similar characteristics between them, list how B is similar to A. Return {num} points as a JSON list of strings. Please focus on specific and granular similarities between the two datasets, your generated characteristic points should apply to all the samples from the two datasets.
Samples from A:
{A}
Samples from B:
{B}
Listing 10: Rubric compilation prompt used in sentiment analysis (differences)
You are a world class data analyst on financial news headlines. You will be given some financial news headline samples from dataset A and dataset B. Based on the provided similar characteristics between them, list how B is {feedback} A. Please focus on granular differences between the two datasets, your generated characteristic points should apply to all the samples from the corresponding dataset (A or B). Return {num} points as a JSON list of strings.
Similar characteristics between A and B:
{similar_points}
Samples from A:
{A}
Samples from B:
{B}
Listing 11: Rubric compilation prompt used in Text2SQL (commonalities)
You are a world class data analyst on database queries in natural language. Given samples from dataset A and dataset B, list how B is {feedback} A. Return {num} points as a JSON list of strings. Please focus on specific and granular similarities between the two datasets, your generated characteristic points should apply to all the samples.
Question samples from A:
{A}
Question samples from B:
{B}
Listing 12: Rubric compilation prompt used in Text2SQL (differences)
You are a world class data analyst on database queries in natural language. Given query samples from dataset A and dataset B. Based on the provided similar characteristics between them, list how B is {feedback} A. Return {num} points as a JSON list of strings. Please focus on specific and granular differences between the two datasets, your generated characteristic points should apply to all the samples from the corresponding dataset (A or B).
Similar characteristics between A and B:
{similar_points}
Question samples from A:
{A}
Question samples from B:
{B}
Listing 13: Rubric compilation prompt used in image classification (commonalities)
Below are {num_image} images from dataset B:
{B}
Below are {num_images} images from dataset A:
{A}
Given some samples from image classification dataset A and dataset B, list how dataset B is similar to dataset A. Return \${num_points} points that summarize the similar characteristics of the two datasets. Focus on the characteristics of the image in terms of how they are structured, styled, or captured (e.g., lighting, background, composition, etc.) rather than the image specifications such as resolution, size, etc. Your generated characteristic points should apply to all the samples from the corresponding dataset (A or B). Output should be a JSON list of strings.
Your listed points:
Listing 14: Rubric compilation prompt used in image classification (differences)
Below are {num_image} images from dataset B:
{B}
Below are {num_images} images from dataset A:
{A}
Given some samples from image classification dataset A and dataset B, list how dataset B is different from dataset A. Similar characteristics are provided below for reference. Return \${num_points} points that summarize the characteristics of the two datasets (e.g., dataset A is ... dataset B is ...). Focus on the characteristics of the images in terms of how they are structured, styled, or captured (e.g., lighting, background, composition, etc.) rather than the image specifications such as resolution, size, etc. Your generated characteristic points should apply to all the samples from the corresponding dataset (A or B). Output should be a JSON list of strings.
Similar characteristics:
\${similar_characteristics}
Your listed points:
Listing 15: Rubric compilation prompt used in web navigation (commonalities)
Given some samples of web navigation tasks and the web accessibility tree of dataset A and dataset B, list how B is {feedback} A. Return {num} points that summarize the characteristics of the two datasets. The two accessibility trees generated from the same website are provided. Please only list out characteristics that are related to the proposed web navigation tasks, the accessibility trees are provided to only help you understand the context of the proposed tasks, therefore do not mention the accessibility tree in your response. Please focus on granular and specific characteristics, and your generated characteristic points should apply to all the samples. Output should be a JSON list of strings.
Accessibility Tree for dataset A:
{A_tree}
Sampled web navigation tasks from dataset A:
{A}
Accessibility Tree for dataset B:
{B_tree}
Sampled web navigation tasks from dataset B:
{B}
Listing 16: Rubric compilation prompt used in web navigation (differences)
Given some samples of proposed web navigation tasks and the web accessibility tree of dataset A and dataset B. Based on the similar characteristics between them. List how B is {feedback} A. Return {num} points that summarize the characteristics of the two datasets. The two accessibility trees of the same website are provided. Please only list out characteristics that are related to the proposed web navigation tasks, the accessibility trees are only provided to help you understand the context of the proposed tasks. Therefore do not list any characteristics that are related to the accessibility tree in your response. Please focus on granular and specific characteristics, and your generated characteristic points should apply to all samples in corresponding dataset (A or B). Output should be a JSON list of strings.
Similar characteristics between A and B:
{similar_points}
Accessibility Tree for dataset A:
{A_tree}
Sampled web navigation tasks from dataset A:
{A}
Accessibility Tree for dataset B:
{B_tree}
Sampled web navigation tasks from dataset B:
{B}
C.5.4 Lens Scoring Prompts
Listing 17: Scorer prompt used in sentiment analysis
You are given similarities and differences between two financial news headline datasets A and B.
Your task is to judge how likely is the given financial news headline comes from dataset {prediction}. Answer your judgement with one of the following strings: "very unlikely", "unlikely", "unsure", "likely", and "very likely".
Similar characteristics between dataset A and B:
{similar_characteristics}
Differences between dataset A and B:
{differences}
Financial news headline sample to be judged:
{example}
Your judgement in JSON format:
Listing 18: Scorer prompt used in Text2SQL
You are given similarities and differences between datasets A and B about database queries in natural language.
Your task is to judge how likely is the given database query in natural language comes from dataset {prediction}. Answer your judgement with one of the following strings: "very unlikely", "unlikely", "unsure", "likely", and "very likely".
Similar characteristics between dataset A and B:
{similar_characteristics}
Differences between dataset A and B:
{differences}
Natural language database query to be judged:
{example}
Your judgement in JSON format:
Listing 19: Scorer prompt used in image classification
You are given similarities and differences between datasets A and B.
Your task is to judge how likely is the given image comes from dataset {prediction}. Answer your judgement with one of the following strings: "very unlikely", "unlikely", "unsure", "likely", and "very likely".
Format:
{format_instructions}
Similar characteristics between dataset A and B:
{similar_characteristics}
Differences between dataset A and B:
{differences}
Image to be judged:
{image}
your judgement in JSON format:
Listing 20: Scorer prompt used in web navigation
You are given similar and different characteristics between two datasets A and B consisting of web navigation tasks.
Your objective is to judge how likely is the given web browsing task comes from dataset {prediction}. Answer your judgement with one of the following strings: "very unlikely", "unlikely", "unsure", "likely", and "very likely".
Format:
{format_instructions}
Similar characteristics between dataset A and B:
{similar_characteristics}
Different characteristics between dataset A and B:
{differences}
Web navigation task to be judged:
{example}
Your judgement in JSON format: